Quadratic functions
I'm confused on how to set up and solve this problem. Among all pairs of numbers with a sum of 205, find the pair whose product is maximum.

Answer by checkley77(12844)

Normally it would be (205/2)^2
But 205/2=102.5^2=10,506.25 ans.
Unless you're looking for awhole number, then:
103*102=10,506 ans.